OpenStudy (anonymous):

If sin (theta) = 1/2 , which could not be the value of theta? a. 30 degrees b. 210 degrees c. 150 degrees d. 390 degrees

OpenStudy (anonymous):

This is how you do it. sin theta = 1/2 theta = sin-1 (1/2) = 30 theta = 30, (180-30)=150, (360+30)=390 therefore answer: B.

OpenStudy (anonymous):

Well that is what I got but I am not Sure. Sorry if I get it wrong

OpenStudy (er.mohd.amir):

principal value is 30

OpenStudy (anonymous):

so it is 210?

OpenStudy (anonymous):

I thinks so, Uhh Yea..

OpenStudy (anonymous):

alright thanks !

OpenStudy (anonymous):

You are very welcome

Nnesha (nnesha):

it CAN'T BE 210 frsure

OpenStudy (anonymous):

What do you think it is @Nnesha ?

Nnesha (nnesha):

sin is positive so either it's in first or 2nd quadrant  cast rule!

Nnesha (nnesha):

 and given sin value is positive so it should be in 1st and 2nd quadrant